# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ 3. Proof sketch for the converse direction of Theorem 4
A graph-nonedge pair $(G,f)$ is a winged graph pair if $G\cup f$ has no $f$-preserving forbidden minor and has an $f$-winged graph minor. The above discussion demonstrates that $(G_{\star},f_{\star})$ is a winged graph pair. The vertex set of any $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$ is written as an indexed set $\{[a_{i}]\}$, where $[a_{i}]$ is $w_{i}$ in one of the graphs in Figure 8. Letting $f_{\star}=uv$, we assume that $u\in[a_{1}]^{-1}$ and $v\in[a_{2}]^{-1}$. Proposition 2, stated below and proved in Appendix B.2, allows us to assume that $[u]^{-1}=\{u\}$ and $[v]^{-1}=\{v\}$, in which case we say that $[G_{\star}\cup f_{\star}]$ has singleton wing tips. See Figure 11(b).
Proposition 2 (Winged graph minors with singelton wing tips).
Consider a winged graph pair $(G,f)$ such that $G\cup f$ is $2$-connected. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then either $(G,f)$ has the $3$-SIP or some $f$-winged graph minor of $G\cup f$ has singleton wing tips.
For any graph $G$, subset $U\subseteq V(G)$, and vertices $s,t\in V(G)\setminus U$, $U$ is an $st$-separator of $G$ if $s$ and $t$ are not contained in the same connected component of $G\setminus U$. An $st$-separator of $G$ is minimal if none of its proper subsets is an $st$-separator of $G$. A pair of vertices $xy$ of $G_{\star}$ is a pin of $[G_{\star}\cup f_{\star}]$ if $\{x,y\}$ is a minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ such that either $x\in[a_{3}]^{-1}$ and $y\in[a_{4}]^{-1}$ or $x\in[a_{5}]^{-1}$ and $y\in[a_{6}]^{-1}$. Note that the paths in $G_{\star}\setminus[f_{\star}]^{-1}$ implied by the existence of $[G_{\star}\cup f_{\star}]$ show that any minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ that is contained in $V(G_{\star})\setminus\{[u]^{-1},[v]^{-1}\}$ and has size at most two is a pin of $[G_{\star}\cup f_{\star}]$. Proposition 3, stated below and proved in Appendix B.3, shows that $[G_{\star}\cup f_{\star}]$ has pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$. See Figure 11(c).
Proposition 3 (Pins of a winged graph minor).
For any winged graph pair $(G,f)$ and any $f$-winged graph minor $[G\cup f]$ whose vertex set is $\{[a_{i}]\}$, there exist pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$.
Next, Proposition 4, stated below and proved in Appendix B.4, allows us to assume that each pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, in which case we call $[G_{\star}\cup f_{\star}]$ chained. See Figure 11(d).
Proposition 4 (Chained winged-graph minors).
For any pin $xy$ of $[G_{\star}\cup f_{\star}]$, observe that there exist unique $\{x,y\}$-components $H_{u}(xy)$ and $H_{v}(xy)$ of $G_{\star}\setminus[f_{\star}]^{-1}$ that contain $u$ and $v$, respectively. Since every pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, each pin of $[G_{\star}\cup f_{\star}]$ other than $xy$ is contained in either $H_{u}(xy)$ or $H_{v}(xy)$, but not both. Hence, we define the ordered sets of pins $\{x_{i}y_{i}\}$ of $[G_{\star}\cup f_{\star}]$ to be the set of all pins of $[G_{\star}\cup f_{\star}]$ indexed such that $H_{v}(x_{i}y_{i})$ contains $x_{i+1}y_{i+1}$. Note that this set contains at least two pins. Proposition 5, stated below and proved in Appendix B.5, shows that $(G_{\star},\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. See Figure 11(d).
Proposition 5 ($(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected).
We are now ready to prove of the converse direction of Theorem 4.
Proof of the converse direction of Theorem 4.
Let $(G,f)$ be a graph-nonedge pair such that no atom of $G\cup f$ that contains $f$ has an $f$-preserving forbidden minor. We will show that $(G,f)$ has the $3$-SIP by induction of the number $n$ of vertices in $G$. When $n\leq 5$, our assumption implies that $G\cup f$ has no forbidden minor, and so Theorem 3 shows that $(G,f)$ has the $3$-SIP. Next, assume that the converse direction of Theorem 4 is true when $n\leq k$, for any integer $k\geq 5$, and we will prove it when $n=k+1$. If $G\cup f$ has no forbidden minor, then Theorem 3 shows that $(G,f)$ has the $3$-SIP. Otherwise, Corollary 3 allows us to assume wlog that $G\cup f$ is an atom. Since atoms are $2$-connected, Proposition 1 shows that $(G,f)$ is a winged graph pair. Consequently, Propositions 2, 3, and 4 allow us to assume that $(G,f)$ has a chained $f$-winged graph minor $[G\cup f]$ with singleton wing tips. Hence, letting $\{x_{i}y_{i}\}$ be the ordered set of pins of $[G\cup f]$, Proposition 5 shows that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected, where $f=uv$. Therefore, by Lemma 2, it suffices to show that $(G\cup\{ux_{2},uy_{2}\},f)$ has the $3$-SIP.
We complete the proof by showing that $(G\cup\{ux_{2},uy_{2}\},f)$ has the $3$-SIP. Observe that $\{u,x_{2},y_{2}\}$ is a clique $wv$-separator of $G\cup\{f,ux_{2},uy_{2}\}$, for some vertex $w\in\{x_{1},y_{1}\}$. Hence, by Lemma 3, it suffices to consider the $\{u,x_{2},y_{2}\}$-component $H$ of $G\cup\{f,ux_{2},uy_{2}\}$ that contains $f$ and show that $(H\setminus f,f)$ has the $3$-SIP. If $(H\setminus f,f)$ does not have the $3$-SIP, then, since $|V(H)|<|V(G)|$, the inductive hypothesis applies and states that $H$ has an $f$-preserving forbidden minor. However, since both $\{x_{1},y_{1}\}$ and $\{x_{2},y_{2}\}$ are minimal $uv$-separators of $G$, there exist paths in the $\{u,x_{2},y_{2}\}$-component of $G\cup f$ that contains $w$ between $w$ and each vertex in the set $\{u,x_{2},y_{2}\}$ such that each path contains exactly one vertex in $\{u,x_{2},y_{2}\}$. This implies that some minor $[G\cup f]_{A}$ is $H$ such that $[f]_{A}$ is $f$. Consequently, $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus f,f)$ has the $3$-SIP. This completes the proof. ∎
4. Proof sketch for the forward direction of main Theorem 4
Given a graph-nonedge pair $(G_{\star},f_{\star})$, we show that if some atom of $G_{\star}\cup f_{\star}$ that contains $f_{\star}$ has an $f_{\star}$-preserving forbidden minor, then $(G_{\star},f_{\star})$ does not have the $3$-SIP. First, we classify the edges of $G_{\star}$. For any graph-nonedge pair $(G,f)$, we treat any edge of $G$ as belonging to $G\cup f$ and $(G,f)$.
Definition 2 (Types of edges of a graph-nonedge pair).
Given a graph-nonedge pair $(G,f)$, the edges of $G$ can be partitioned into the following four classes. Let $[G\cup f]$ be the minor by contracting any one edge of $G$. Then, the contracted edge is of Type
  1. (1)
    if $[G\cup f]$ has no $[f]$-preserving forbidden minor,
  2. (2)
    if $[G\cup f]$ some an $[f]$-preserving forbidden minor, but no atom of $[G\cup f]$ that contains $[f]$ has such a minor,
  3. (3)
  4. (4)
    otherwise, and in which case the edge is called reducing.
Definition 3 (Minimal pairs).
A graph-nonedge pair $(G,f)$ is minimal if $G\cup f$ is an atom that has an $f$-preserving forbidden minor and no reducing edge.
4.0.1. Example
A graph-nonedge pair $(G,f)$ is not minimal if it is the pair (i) on the right in Figure 3 since $G\cup f$ has no $f$-preserving forbidden minor, (ii) in Figure 4(a) even though $G\cup f$ has an $f$-preserving forbidden minor, since $G\cup f$ is not an atom, or (iii) on the bottom-right of Figure 4(c) even though $G\cup f$ is an atom that has an $f$-preserving forbidden minor, since the edge whose contraction yields an $f$-retaining $K_{5}$ minor is reducing. $(G,f)$ is minimal if it is any of the pairs in Figure 12 or if $G\cup f$ is any of the graphs in Figures 9(a), 9(b), 9(d), 9(f), 9(g), or 9(i).
Lemmas 1 and 3 allow us to assume wlog that $(G_{\star},f_{\star})$ is minimal. Hence, the forward direction of Theorem 4 follows from Theorem 7, below. A squared edge-length map $\ell$ is a proper non-$d$-SIP map for a graph-nonedge pair $(G,f)$ if $\ell$ maps each edge of $G$ to a positive number and the CCS $\Omega^{3}_{f}(G,\ell)$ contains exactly two distinct positive real numbers. For example, the squared edge-length map in Figure 1(a) is a proper non-$1$-SIP map for the graph-nonedge pair in this figure, as illustrated in Example 1.0.1.
Theorem 7 (Every minimal pair has a proper non-$3$-SIP map).
We spend the remainder of the section proving Theorem 7. The proof is by induction on the number $n$ of vertices in $G_{\star}$. Since $G_{\star}\cup f_{\star}$ has a forbidden minor, we have $n\geq 5$. When $n=5$, $G_{\star}\cup f_{\star}$ is $K_{5}$, and so Proposition 6, stated below and proved in Appendix C.1, shows that $(G_{\star},f_{\star})$ has a proper non-$3$-SIP map.
Figure 12. All minimal graph-nonedge pairs with no Type (2) edge. See Proposition 6.
Proposition 6 (No Type (2) edge implies non-$3$-SIP map).
In the inductive step, Proposition 6 allows us to assume wlog that $G_{\star}$ has a Type (2) edge. We now present several definitions that will allow us to exploit the presence of such an edge. Consider a graph $G$, any of its edges $e$, any of its induced subgraphs $H$ and $H^{\prime}$, a minor $[G]_{A}$, and a minor $[H]_{B}$. $H\setminus H^{\prime}$ is the subgraph of $G$ induced by $V(H)\setminus V(H^{\prime})$. $[G]_{A}$ is an extension of $[H]_{B}$ if $|V([G]_{A})|=|V([H]_{B})|$ and, for any two vertices $u,v\in V(H)$, $[u]_{A}=[v]_{A}$ if and only if $[u]_{B}=[v]_{B}$. If $H$ contains $e$, then $e$ being contracted, preserved, or retained in $[H]_{B}$ are well-defined concepts (see Section 2). Otherwise, $e$ is contracted in $[H]_{B}$ if it is contracted in every extension of $[H]_{B}$ and preserved (resp. retained) in $[H]_{B}$ if it is preserved (resp. retained) in at least one extension of $[H]_{B}$. $[G]_{A}$ is $e$-separating if it has an $[e]_{A}$-preserving forbidden minor but none of its atoms that contain $[e]_{A}$ has such a minor. For the following, refer to Figure 7. $[H]_{A}$ is the subgraph of $[G]_{A}$ induced by $[V(H)]$. For any induced subgraph $J$ of $[G]_{A}$, $J^{-1}$ is the subgraph of $G$ induced by $\bigcup_{u\in J}u^{-1}$. Note that $[H]_{A}^{-1}$ is not necessarily $H$.
4.0.2. Example
Let $G\cup f$ be the graph in Figure 9(d), where $f$ is the green edge, and let $[G\cup f]$ be its minor shown in Figure 9(e). Observe that the subgraph $H$ of $G\cup f$ consisting of the bottom five vertices and the vertex incident on the two blue edges has an $f$-preserving $K_{5}$ minor. $[G\cup f]$ is an $f$-separating minor since its $K_{5}$ subgraph is its only atom that has an $f$-preserving forbidden minor. Lastly, letting $J$ be the subgraph of $G\cup f$ induced by the bottom five vertices, note that $[J]^{-1}=H$.
Definition 4 (Expanded edge-separating pairs and cms’s).
Let $G$ be a graph and $e$ be any of its edges. An $e$-separating pair $(C,M)$ of an $e$-separating minor $[G]$ contains an atom $M$ of $[G]$, which has an $[e]$-preserving forbidden minor, and a minimal clique $[ux]$-separator $C$ of $[G]$, for some endpoint $[u]$ of $[e]$ and any vertex $[x]$ of $M\setminus C$. Consider the $C$-components $I$ and $H$ of $[G]$ that contain $M$ and $[e]$, respectively. $C$, $I$, and $H$ are the cms, minor-component, and $e$-component of $(C,M)$, respectively.
Next, the pair $(C^{-1},M^{-1})$ is an expanded $e$-separating pair of $G$, and $C^{-1}$, $I^{-1}$, and $H^{-1}$ are the expanded cms, expanded minor-component, and expanded $e$-component of $(C^{-1},M^{-1})$, respectively. $(C^{-1},M^{-1})$ is a top-level expanded $e$-separating pair of $G$ if no proper subgraph of $H^{-1}$ is the expanded $e$-component of any expanded $e$-separating pair of $G$. $(C,M)$ is a top-level $e$-separating pair of $[G]$ if $(C^{-1},M^{-1})$ is in the top-level of $G$.
Finally, an $e$-separating cms $E$ of $[G]$ is a cms of some $e$-separating pair $(E,M)$ of $[G]$. We treat the minor-component and $e$-component of $(E,M)$ as belonging to $E$. $E^{-1}$ is an expanded $e$-separating cms of $G$, and is in the top-level if $(E^{-1},M^{-1})$ is in the top-level.
4.0.3. Example
Let $G\cup f$ be the graph in Figure 9(b), where $f$ is the green edge, $[G\cup f]$ be the minor shown in Figure 9(c), and $C$ and $M$ be the purple and $K_{5}$ subgraphs of $[G\cup f]$, respectively. Observe that $(C,M)$ is an $f$-separating pair of $[G\cup f]$, and hence $C$ is an $f$-separating cms. The $f$-component $H$ of $C$ is the subgraph in the dashed black box and the minor component is $M$. The subgraph $C^{-1}$ of $G\cup f$ is a top-level expanded $f$-separating cms whose expanded $f$-component and expanded minor component are $H^{-1}$ and $M^{-1}$, respectively. Note that the orange subgraph of $G\cup f$ is also a top-level expanded $f$-separating cms whose expanded $f$-component and expanded minor component are $H^{-1}$ and $M^{-1}$, respectiveely.
Finally, the above statements are true for the pairs of graphs in Figures 9(g) and 9(h) and in Figures 9(i) and 9(j), except, in the latter case, $E^{-1}$ is the connected subgraph consisting of one purple edge and one blue edge.
By definition, contracting any Type (2) edge of $G_{\star}\cup f_{\star}$ yields an $f_{\star}$-separating minor. Hence, Proposition 7, stated below and proved in Appendix C.2, shows that $G_{\star}\cup f_{\star}$ has a top-level expanded $f_{\star}$-separating cms.
Next, let $C^{-1}$ be any top-level expanded $f_{\star}$-separating cms of $G_{\star}\cup f_{\star}$, $H^{-1}$ be its expanded $f_{\star}$-component, and $G^{\prime}=G_{\star}\setminus(H^{-1}\setminus C^{-1})$. Using the inductive hypothesis and Proposition 8, stated below and proved in Appendix C.3, it is easy to show that some nonedge $f^{\prime}$ of $C^{-1}$ is such that $(G^{\prime},f^{\prime})$ has a proper non-$3$-SIP map $\ell^{\prime}$.
Proposition 8 (Minimal pair inductive structure).
The goal is now to use $\ell^{\prime}$ to construct a proper non-$3$-SIP map $\ell$ for $(G_{\star},f_{\star})$. For example, if $G_{\star}\cup f_{\star}$ is the graph in Figure 9(b) (resp. 9(g)), where $f_{\star}$ is the green edge, then we can take $G^{\prime}\cup f^{\prime}$ to be the graph in Figure 9(a) (resp. 9(f)), where $f^{\prime}$ is the green edge. It is not hard to see that $\ell$ can be chosen to agree with $\ell^{\prime}$ on edges of $G^{\prime}$ and to assign squared lengths for the two remaining edges of $G_{\star}$ whose sum and difference are the two values in the CCS $\Omega^{3}_{f^{\prime}}(G^{\prime},\ell^{\prime})$. This ensures that $\ell$ has the desired properties. Similarly, $\ell$ can be extended to obtain a non-$3$-SIP map for the pair in Figure 9(d) (resp. 9(i)). To implement the above strategy in general, Proposition 9, stated below and proved in Appendix C.4, lists all possible graphs for $H^{-1}$. Combining this with Proposition 10, stated below and proved in Appendix C.5, completes the proof by induction.
Proposition 9 (Pruned possible graphs for top-level $H^{-1}$).
Let $H^{-1}$ be the expanded $f$-component of any top-level expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair. Then, $H^{-1}$ is one of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
Proposition 10 (Extending non-$3$-SIP maps for minimal pairs).
Let $G^{\prime}$ be any of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, let $H$ be the subgraph of $G$ induced by the set $\{w_{i}\}$, let $f$ be the red edge, and let $f^{\prime}$ be the nonedge of $H$ that does not share an endpoint with $f$ if possible, or any other nonedge of $H$ otherwise. Also, let $G=(G^{\prime}\setminus f)\cup f^{\prime}$. Lastly, let $\ell_{1}$ and $\ell_{2}$ be squared edge-length maps for $H$ that assign each edge a positive real number and differ only in the numbers assigned to $f^{\prime}$. If the linkages $(H,\ell_{1})$ and $(H,\ell_{2})$ both have at least one $3$-realization, then there exist squared edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for $G$ such that
  1. (i)
  2. (ii)
    $\ell^{\prime}_{i}$ agrees with $\ell_{i}$ on the edges of $H$, and
  3. (iii)
Before proving Theorem 7 and the forward direction of Theorem 4, we sketch the the proofs of Propositions 8 and 9, as they are important and highly technical. Lemmas 4 and 5, stated below and proved in Appendix C.3.1, describe the combinatorial structure of $C^{-1}$ and $H^{-1}$, respectively. For any induced subgraphs $H_{1}$ and $H_{2}$ of $G$, $H_{1}\cup H_{2}$ is the subgraph of $G$ induced by $V(H_{1})\cup V(H_{2})$.
Figure 13. Top-level expanded $f$-separating cms’s $C^{-1}$ of minimal graph-nonedge pairs. See Definition 4 and Lemma 4.
Figure 14. Graphs in Lemma 4. Dashed line-segments may or may not be edges.
Lemma 4 (Unpruned possible graphs for top-level $C^{-1}$).
  1. (1)
    $C^{-1}$ is one of the graphs in Figure 13 and
  2. (2)
    the minor $[H^{-1}\cup I^{-1}]$ obtained by contracting every edge in both $H^{-1}\setminus C^{-1}$ and $I^{-1}\setminus C^{-1}$ is one of the graphs in Figure 14 such that $[C^{-1}]$ is the subgraph induced by the set $\{w_{i}\}$.
Lemma 5 (Structure of top-level $H^{-1}$).
These two lemmas are used to prove Lemma 6, stated below and proved in Appendix C.3.2, which eliminates the graphs in Figures 13(e) - 13(i) as possibilities for $C^{-1}$.
Lemma 6 (Pruned possible graphs for $C^{-1}$).
If $C^{-1}$ is a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, then $C^{-1}$ is one of the graphs in Figures 13(a)-13(d).
These three lemmas are used to prove Lemma 7, stated below and proved in Appendix C.3.3, which gives the unpruned set of all possible graphs for $H^{-1}$. The proof of this lemma is highly technical and has many cases.
Figure 15. See Lemma 7. All possible expanded $f$-components of an expanded $f$-separating cms that has fewer than five vertices and is in the top-level of a minimal graph-nonedge pair $(G,f)$. Dashed line-segments may be edges or nonedges. Hollow vertices may or may not be present. In 15(h), either $u$ neighbors $w_{1}$ or at least one of the hollow vertices is present. In 15(e), 15(i), or 15(l), either $u$ neighbors $w_{3}$ or at least one of the hollow vertices is present. In 15(k), either $u$ neighbors some vertex in $\{w_{3},w_{4}\}$ or at least one of the hollow vertices is present.
Lemma 7 (Unpruned possible graphs for top-level $H^{-1}$).
If $C^{-1}$ is a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, then the expanded $f$-component of $C^{-1}$ is one of the graphs in Figure 15 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$.
After reducing the possible graphs for $H^{-1}$ to those in Figures 15(a) - 15(m) (Lemma 44 in Appendix C.3.4), proving Proposition 8 is just a matter of case checking (Lemma 47 in Appendix C.3.4). Finally, Proposition 9 is proved by eliminating the graphs in Figures 15(f) - 15(m) as cases for $H^{-1}$ (Lemmas 51 - 54 in Appendix C.4).
Proof of Theorem 7.
We show that a minimal graph-nonedge pair $(G,f)$ has a proper non-$3$-SIP map by induction on the number $n$ of vertices in $G$. Since $G\cup f$ has a forbidden minor, we have $n\geq 5$. When $n=5$, $G\cup f$ is $K_{5}$, and so the theorem follows from Proposition 6. Assume that the theorem is true when $n\leq k$, for any integer $k\geq 5$, and we will show that it is true when $n=k+1$. The same proposition allows us to assume that $(G,f)$ has a Type (2) edge. Hence, $(G,f)$ has a top-level expanded $f$-separating cms $C^{-1}$, by Proposition 7. Let $H^{-1}$ be the expanded $f$-component of $C^{-1}$ and $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Consider any minimal graph-nonedge pair $(J,f^{\prime})$ obtained by applying Proposition 8 to $(G,f)$ and $C^{-1}$. Proposition 9 tells us $H^{-1}$ is one of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$. Hence, Proposition 8 allows us to choose $f^{\prime}=w_{1}w_{3}$ if $H^{-1}$ is the graph in Figure 15(c), and $f^{\prime}=w_{2}w_{4}$ if $H^{-1}$ is the graph in Figure 15(d).
Proof of the forward direction of Theorem 4.
Let $(G,f)$ be a graph-nonedge pair such that some atom of $G\cup f$ that contains $f$ has an $f$-preserving forbidden minor. Corollary 3 allows us to assume wlog that $G\cup f$ is an atom. If $(G,f)$ has a reducing edge, then let $[G\cup f]$ be the minor obtained by contracting any such edge. By definition, $[G\cup f]$ is $f$-retaining and has an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. Hence, by Lemma 1, if $([G\cup f]\setminus[f],[f])$ does not have the $3$-SIP, then neither does $(G,f)$. Therefore, we can assume wlog that $(G,f)$ has no reducing edge. Combining this with the fact that $G\cup f$ is an atom shows that $(G,f)$ is minimal, and thus the theorem follows from Theorem 7. ∎
5. Open problems and conjectures
Open Problem 1.
There is likely such an algorithm based on the polynomial time of [17] for finite forbidden minors.
Open Problem 2.
Use the characterization in Theorem 4 to characterize $3$-convex graph-nonege-set pairs, similar to how the characterization of $2$-convex graph-nonege-set pairs in [11] is based on their characterization of graph-nonege pairs with the $2$-SIP.
We believe our characterization in Theorem 4 extends to any dimension as follows.
Conjecture 1.
A graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbidden minor.
Since the number of forbidden minors for $d$-flattenability is conjectured to grow asymptotically with the number of forbidden minors for partial $d$-trees [5], which is at least $75$ for $d=4$, a proof of this conjecture must avoid using properties of the explicit $d$-flattenability forbidden minors. To this end, we posit the following problem.
Open Problem 3.
For any integer $d\geq 0$, give a characterization of the set of $d$-flattenability forbidden minor, other than the set of minors that characterize $d$-flattenable graphs.
Appendix A Proofs of the lemmas in Section 2
A.1. Proof of Lemma 2 [Edge deletions that preserve $d$-Cayley-connectivity]
Lemma 2 is proved using Lemma 8, below. For any integer $n>0$, consider a subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections. For any point $x\in S^{\prime}$, the fiber $S(x)$ of $x$ through $S$ is the preimage of $x$ under the projection map from $S$ to $S^{\prime}$. Recall that in this paper, all sets are closed and compact, all maps are closed, and connected implies path-connected, unless otherwise specified.

### v2

@@ 4   Proof of the converse direction of the main Theorem 4
A graph-nonedge pair $(G,f)$ is a winged graph pair if $G\cup f$ has no $f$-preserving forbidden minor and has an $f$-winged graph minor. The above discussion demonstrates that $(G_{\star},f_{\star})$ is a winged graph pair. Let $f_{\star}=uv$. The vertex set of any $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$ is written as an indexed set $\{[a_{i}]\}$, where $u\in[a_{1}]^{-1}$, $v\in[a_{2}]^{-1}$, $[a_{3}]$ and $[a_{4}]$ are neighbors of $[a_{1}]$, and $[a_{5}]$ and $[a_{6}]$ are neighbors of $[a_{2}]$. Proposition 2, below, allows us to assume that $[a_{1}]^{-1}=\{u\}$ and $[a_{2}]^{-1}=\{v\}$, in which case we say that $[G_{\star}\cup f_{\star}]$ has singleton wing tips. See Figure 12(b).
Proposition 2 is proved using Lemma 6, below. For any $f$-winged graph minor $[G\cup f]$ of a winged graph pair $(G,f)$, with $f=uv$, we call vertices $s\in[u]^{-1}$ and $t\in[v]^{-1}$ wing tips of $[G\cup f]$ if either they are the only vertices in $[u]^{-1}\cup[v]^{-1}$ or they separate every other vertex in $[u]^{-1}\cup[v]^{-1}$ from every vertex not in $[u]^{-1}\cup[v]^{-1}$. Consider the following operation, similar to a vertex exchange, that transforms one minor $[G]_{1}$ of a graph $G$ into another minor $[G]_{2}$. Refer to Figure 13. Consider an edge $[ab]_{1}$ of $[G]_{1}$, the subgraph $H$ of $G$ induced by the set $[a]_{1}^{-1}$, a connected subgraph $J$ of $H$, and a connected component $K$ of $H\setminus J$ that contains a vertex with a neighbor in the set $[b]_{1}^{-1}$. The minor $[G]_{2}$ obtained by replacing $[a]_{1}^{-1}$ with $[a]_{1}^{-1}\setminus V(K)$ and $[b]_{1}^{-1}$ with $[b]_{1}^{-1}\cup V(K)$ is said to be obtained via the component exchange from $[a]_{1}^{-1}$ to $[b]_{1}^{-1}$ that exchanges $K$.
Lemma 6 (Wing tips of winged graph minors).
For any winged graph pair $(G,f)$, with $f=uv$, and any $f$-winged graph minor $[G\cup f]$, there exists an $f$-winged graph minor $[G\cup f]_{1}$ with wing tips such that $[u]_{1}^{-1}\cup[v]_{1}^{-1}$ is a subset of $[u]^{-1}\cup[v]^{-1}$ and both $[f]_{1}^{-1}$ and $[f]^{-1}$ contain the same edges of $G$.
Proof.
Let $[G\cup f]_{1}$ be any $f$-winged graph minor whose vertex set is $\{[a_{i}]_{1}\}$ such that (i) $[u]_{1}^{-1}\cup[v]_{1}^{-1}$ is a subset of $[u]^{-1}\cup[v]^{-1}$ and both $[f]_{1}^{-1}$ and $[f]^{-1}$ contain the same edges of $G$ and (ii) $[u]_{1}^{-1}\cup[v]_{1}^{-1}$ has minimum size over all such minors satisfying (i). Assume there exist at least two vertices in $[u]_{1}^{-1}$ with neighbors in $[a_{3}]_{1}^{-1}\cup[a_{4}]_{1}^{-1}$. Then, we can choose $x$ to be one of these vertices other than $u$. Let $H$ be the subgraph of $G\cup f$ induced by $[u]_{1}^{-1}$ and let $J$ be the connected component of $H\setminus x$ that contains $u$. Assume wlog that $x$ has a neighbor in $[a_{3}]_{1}^{-1}$. If $J$ contains some vertex with a neighbor in $[a_{4}]_{1}^{-1}$, then let $[G\cup f]_{2}$ be the minor obtained via the vertex exchange from $[u]_{1}^{-1}$ to $[a_{3}]_{1}^{-1}$ that exchanges $x$ and fixes $J$. Else, if $J$ contains some vertex with a neighbor in $[a_{3}]_{1}^{-1}$ and $x$ has a neighbor in $[a_{4}]_{1}^{-1}$, then let $[G\cup f]_{2}$ be the minor obtained via the vertex exchange from $[u]_{1}^{-1}$ to $[a_{4}]_{1}^{-1}$ that exchanges $x$ and fixes $J$. Lastly, if no vertex in $[a_{4}]_{1}^{-1}$ neighbors either $x$ or any vertex in $J$, then some connected component $K$ of $H\setminus x$ other than $J$ contains some vertex with a neighbor in $[a_{4}]_{1}^{-1}$. In this case, let $[G\cup f]_{2}$ be the minor obtained via the component exchange from $[u]_{1}^{-1}$ to $[a_{4}]_{1}^{-1}$ that exchanges $K$.
Next, in each case above, the neighborhoods of vertices in $[G\cup f]_{1}$ along with Lemma 5 show that $[G\cup f]_{2}$ is an $f$-winged graph minor such that $[u]_{2}^{-1}\cup[v]_{2}^{-1}$ is a proper subset of $[u]_{1}^{-1}\cup[v]_{1}^{-1}$ and both $[f]_{2}$ and $[f]_{1}$ contain the same edges of $G$. However, this contradicts Property (ii). Therefore, $[u]_{1}^{-1}$ contains exactly one vertex with a neighbor in $[a_{3}]_{1}^{-1}\cup[a_{4}]_{1}^{-1}$. Symmetrically, $[v]_{1}^{-1}$ contains exactly one vertex with a neighbor in $[a_{5}]_{1}^{-1}\cup[a_{6}]_{1}^{-1}$. Thus, $[G\cup f]_{1}$ satisfies Property (i) and has wing tips, as desired. ∎
Proposition 2 additionally requires tool Lemmas 2 and 3 from Section 2 that are of independent interest and are used several times throughout this paper. We prove them now in the following subsection.
4.1   Proofs of Tool Lemmas 2 and 3
Lemma 2, about edge deletions which preserve $d$-SIP, is proved using Lemma 7, below. For any dimension $n>0$, consider a subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections given by the projection map $\pi$. For any point $x\in S^{\prime}$, we denote the fiber $\pi^{-1}(x)$ by $S(x)$. Recall that in this paper, all sets are closed and compact, all maps are closed, and connected implies path-connected, unless otherwise specified.